FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF THE N_{2} CO_{2} VAN DER WAALS COMPLEX

\underline{EDWIN S. CONTRERAS}, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The microwave spectrum of the weakly-bound complex of nitrogen and carbon dioxide was measured by Fourier transform microwave spectroscopy. A previous determination of the structure of this molecule was performed in a rotationallyresolved infrared study in the region of ν_{3} of CO_{2}^{a}. We have observed only a-type transitions, and have measured the hyperfine splitting due to the two ^{14}N nuclei. Consistent with the geometry proposed by Walsh et. al., the structure of the complex is T-shaped, with the CO_{2} forming the cross of the T.

aM. A. Walsh, T. R. Dyke, and B. J. Howard, J. Mol. Struct. 189, 111, (1988)